OpenStudy (anonymous):

What are the x-intercepts of the quadratic function? y = x2 + x – 2 A. 2 and 1 B. –2 and 1 C. 2 and –1 D. 1 A ?

OpenStudy (anonymous):

it's x^2

OpenStudy (radar):

Solve by inspection. What value is x for y to = 0?

OpenStudy (radar):

Note the answers only involve 1, -2, -1, or 2 start out testing the 1.

OpenStudy (radar):

Don't be shy. I will test the '1'. 1^2 + 1 - 2= 1 + 1 -2 = 2-2=0 y=0 That eflatt13 is a winner, but there is still one more. Can you find it?

OpenStudy (superhelp101):

sorry to bother you radar, but can you help me when you're done. I would appreciate it much! thanks! again sorry to bother you, only question i have, i promise

OpenStudy (radar):

No problem @superhelp101, I may be of some help, but am not infallible.